how many milliliter of a 19% alcohol solution must be added to 50 milliliter of 10% alcohol solution to make a 15% alcohol solution

Found 2 solutions by mananth, robertb:
Answer by mananth(16946)

how many milliliter of a 19% alcohol solution must be added to 50 milliliter of 10% alcohol solution to make a 15% alcohol solution
..
19% alcohol solution : x ml
10% alcohol solution 50 ml
15% alcohol solution : x+50 ml
...
0.19x+0.1*50=0.15(x+50)
0.19x+5=0.15x+7.5
0.19x-0.15x=7.5-5
0.04x = 2.5
x=2.5/0.04
= 62.5 ml of 19%
...
m.ananth@hotmail.ca

Answer by robertb(5830)

Let x = #ml of 19% alcohol solution that must be added. Then
0.19x + 0.10*50 = 0.15(x+50),
0.19x + 5 = 0.15x + 7.5,
0.04x = 2.5,
x = 62.5 ml of the 19% alcohol solution must be added.